# Constraining Statistical Isotropy using 21cm Power Spectrum and Bispectrum

### v1

@@ 1 Introduction
In the present work, our objective is to study both dipolar and quadrupolar asymmetries in the primordial power spectra using 21cm intensity mapping during post reionization era, i.e., $z\lesssim 7$. We put constraints on the anisotropy parameters using power and bispectrum employing Fisher forecasts for intensity mapping surveys: SKA-Mid [18], HIRAX [19, 20] and PUMA [21]. To our knowledge, only two such studies have hitherto been done in the literature. In [22], Shiraishi et. al, tested the statistical isotropy violation using 21cm signals during the cosmic dark ages where the effect of foreground isn’t that prominent. In their work, they focused only on the 21cm power spectrum and gave forecast on both dipolar and quadrupolar anisotropy parameters for the futuristic radio telescope SKA. They found $1\sigma$ error on dipolar modulation amplitude is $\lesssim 10^{-3}$ compared to $\sim 10^{-3}$ for an ideal CMB experiment. For quadrupolar anisotropy, the constraints were $\lesssim 10^{-3}$. In another study done by Li et al. [23], authors tested the dipolar anisotropy from the epoch of reionization for SKA and Omniscope. They constrained the dipolar anisotropy parameter using the 21cm power spectrum while accounting for foregrounds by neglecting the modes with $k_{||}\geq k_{||,\mathrm{min}}$. They found that SKA Phase 2 is capable of constraining the dipolar modulation amplitude parameter $\sim 10^{-2}$ for Fourier modes that range between $0.056$ and $0.15\ \mathrm{Mpc}^{-1}$. They also found that the constraints on the parameter can be improved by an order of magnitude using Omniscope.
@@
The paper is structured in the following manner. In §2, we start with a brief review of the physics of the 21cm line including a discussion of (a) various survey modes of 21cm or HI intensity mapping and (b) foreground problem and how it is avoided. This is followed by a brief discussion of primordial asymmetries considered in this paper in §3. After this, in §4, we present the calculation of power spectrum and bispectrum in the presence of asymmetries. The Fisher forecasts results are presented in §5. The effects of non-linearity decrease as $z$ becomes larger. We find relatively larger contribution to the Fisher matrix at large $z$ values. Thus, for the purpose of this paper, we restrict ourselves to the linear regime. Also, while making forecasts, we consider all relevant effects like redshift space distortions, fingers of God, instrumental noise, effect of telescope beam, etc. We conclude in §6. In this paper, we use fiducial $\Lambda$CDM cosmology with the best fit parameters from Planck 2018 paper [26]: $h$ = 0.6766, $\Omega_{\mathrm{b}}h^{2}=0.02242$ , $\Omega_{\mathrm{c}}h^{2}$ = 0.11933, $n_{\mathrm{s}}$ = 0.9665, $A_{\mathrm{s}}=2.142\times 10^{-9}$.
@@ 2.3 Foreground Problem and its Avoidance
$|k_{\parallel}|>\tan\alpha(z)k_{\perp},$ (2.8)
@@ 3.1 Dipolar Asymmetry
Large-scale observations of the CMB show a deviation from the isotropic distribution, demonstrating evidence of anomalies that conflict with the cosmological principle. As discussed above, the observations of WMAP and Planck indicate statistical anisotropy aka hemispherical power asymmetry in the CMB temperature map [25, 58, 22]. This is modelled as a dipolar modulation in the statistically isotropic temperature field $T^{\mathrm{iso}}$ as $T(\hat{\mathbf{n}})=T^{\mathrm{iso}}(\hat{\mathbf{n}})(1+A\hat{\mathbf{n}}\cdot\hat{\mathbf{p}})$, where $\hat{\mathbf{p}}$ is the preferred direction and $\hat{\mathbf{n}}$ is the observation direction. The asymmetry magnitude $A\sim 0.07$ and the direction in galactic coordinates is $\hat{\mathbf{p}}=(l,b)=(205^{\circ},-20^{\circ})$ [7]. Phenomenologically, this power asymmetry can be studied by introducing a position dependent dipolar anisotropy in the primordial curvature perturbations [59, 22]. Some papers also explain the effect by introducing inhomogeneity in the primordial power spectrum [13, 60]. Given large-scale nature of dipolar anisotropy, it may have connections with inflationary physics. In this regard several, theoretical models have been studied [61, 62, 63]. Some of these models have been ruled out because of their isotropic predictions [64, 65].
@@
where $\hat{\mathbf{p}}$ is the direction of asymmetry, $\hat{\mathbf{n}}$ the vector along the LoS which we choose to be along Z axis for convenience and $P_{\zeta}^{0}(k)$ was defined above. Also, we have taken a scale dependent modulation $A(k)$, motivated by the fact that the hemispherical power asymmetry in CMB is present at large angular scales and typically dies out beyond $\ell\gtrsim 64$ [5]. The scale dependence is also suggested by quasar density, where at low $z$ the dipole modulation vanishes at scales $k\sim 1\ \mathrm{Mpc}^{-1}$ [12]. We can factor out the scale dependence in $A(k)$ and write $A(k)=A_{0}F(k)$, where $A_{0}=2\hat{\mathbf{p}}\cdot\hat{\mathbf{n}}=2\hat{\mathbf{p}}\cdot\hat{z}$, the parameter that we later constrain using the Fisher formalism. With all this, (3.2) can be written as
@@
For the present study, we consider the following scale dependencies $F(k)$
$F(k)=\left(1-\frac{k}{k_{A}}\right)^{n},\ n=0,1,2$ (3.4)
where $k_{A}=1\ \mathrm{Mpc}^{-1}$. The case $n=0$ correspond to no scale dependence that we compare with $n=1,2$ used in [22] to discuss the detectability of 21cm power spectrum. This is also motivated by some theoretical models where varying spectral index $n_{\mathrm{s}}$ can destroy asymmetry at some scale $k_{A}$ [66].
@@ 3.2 Quadrupolar Asymmetry
Although statistical isotropy is the prediction of standard inflationary theory, theoretical research has shown that inflationary models can be constructed that deviate from it. One such model of primordial anisotropy is when the inflaton interacts with a vector field having a nonzero vacuum expectation, transferring the anisotropy [17]. Such a model provides a quadrupolar anisotropic primordial power spectrum. We consider the primordial scalar power spectrum to have a quadrupolar directional dependence about some axis $\hat{\mathbf{d}}$ [58, 25, 67]. Without loss of generality, this axis can be chosen along the Z axis. Thus, we get the following expression
$P_{\zeta}(\mathbf{k})\approx P_{\zeta}^{0}[1+g_{0}g(k)\mathcal{P}_{2}(\hat{k}\cdot\hat{z})]$ (3.5)
@@
here $k_{*}=0.05\mathrm{Mpc}^{-1}$ is the pivot scale chosen by Planck, $\mathcal{P}_{n}(x)$ is the nth order Legendre polynomial and $g_{0}$ is the amplitude of the quadrupolar anisotropy whose magnitude depends on model parameters, particularly in vector field models where it is proportional to the ratio of energy densities [17]. We later constrain this parameter using Fisher formalism.
@@ 4.1 Power Spectrum
where $\bar{T}_{\mathrm{b}}$ is the mean brightness temperature, the term $D_{\mathrm{FoG}}(\mathbf{k})=\text{exp}[-k^{2}\mu^{2}\sigma^{2}_{\mathrm{P}}(z)]$ corresponds to the Fingers of God (FoG) effect, with $\sigma_{\mathrm{P}}$ denoting the LoS dispersion in relative galaxy velocities. In this paper, we use $\sigma_{\mathrm{p}}=350/H(z)$, as per Ref. [71]. Further, $b_{1}$ is the neutral hydrogen bias given by
$b_{1}(z)=1.333+0.1809z+0.05302z^{2}-0.0008822z^{3}$ (4.3)
@@
where $\mathcal{T}(k)$ is the relevant transfer function and 00 in the superscripts are the quantities in the absence of primordial asymmetries. In the presence of asymmetries, this equation gets modified to
@@
$\mathrm{Quadrupolar}$ $:P_{\mathrm{m}}^{0}(k,z)\longrightarrow P_{\mathrm{m}}(\mathbf{k},z)=P_{\mathrm{m}}^{0}(k,z)\Big{[}1+g_{0}g(k)\mathcal{P}_{2}(\mu)\Big{]}$ (4.7)
where the scales dependencies $F(k)$ and $g(k)$ are given in (3.4) and (3.6). Thus our final expressions for the 21cm power spectrum in the presence of the dipolar and quadrupolar anisotropies are simplified to
$P_{21}^{\mathrm{D}}(\mathbf{k},z,\hat{\mathbf{n}})$ $=P^{0}_{21}(\mathbf{k},z)+A_{0}(P^{0}_{21}(\mathbf{k},z)-P_{\mathrm{N}}(\mathbf{k},z))F(k)$ (4.8)
$P_{21}^{\mathrm{Q}}(\mathbf{k},z)$ $=P^{0}_{21}(\mathbf{k},z)+g_{0}(P^{0}_{21}(\mathbf{k},z)-P_{\mathrm{N}}(\mathbf{k},z))g(k)\mathcal{P}_{2}(\mu)$ (4.9)
@@ 4.2 Bispectrum
The Dirac Delta imposes (a) the condition that the vectors $\mathbf{k}_{1}$, $\mathbf{k}_{2}$ and $\mathbf{k}_{3}$ form three sides of the triangle and (b) homogeneity as before. Apparently, there are 9 variables (three components each for $\mathbf{k}_{i}$, $i=1,2,3$) in the bispectrum expression (4.10) but three variables are eliminated by the Dirac Delta condition $\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}=0$, thereby leading to 6 independent variables. Also, the bispectrum is invariant under a rotation in the plane perpendicular to the LoS [72]. So we are left with 5 variables that completely describe the bispectrum. We can choose three variables to define the shape of the triangle and remaining two for describing its orientation with respect to the LoS. Thus finally, 21cm bispectrum, (4.10) can be written as $B_{21}(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3},z)=B_{21}(k_{1},k_{2},k_{3},\mu,\phi,z)$, where $k_{1}$, $k_{2}$ and $k_{3}$ are the sides of the triangle, $\mu$ is the cosine of angle between $\mathbf{k}_{1}$ and Z axis (the chosen LoS). Also, $\phi$ is the azimuthal angle about $\mathbf{k}_{1}$ [68].
@@
here $\mu_{12}=(\mu_{1}k_{1}+\mu_{2}k_{2})/k_{12}$, $k_{12}^{2}=(\mathbf{k}_{1}+\mathbf{k}_{2})^{2}$, $F_{2}(\mathbf{k}_{1},\mathbf{k}_{2})$ $G_{2}(\mathbf{k}_{1},\mathbf{k}_{2})$ are the mode-coupling part of the matter density contrast & matter velocity [74] and $S_{2}(\mathbf{k}_{1},\mathbf{k}_{2})=(\hat{k}_{1}\cdot\hat{k}_{2})^{2}-1/3$ is the tidal kernel [75]. The neutral hydrogen biases $b_{i}(z)$, $i=1,2$ and $b_{s^{2}}$ relate the statistics of observed tracers with the dark matter distribution. The bias $b_{1}$ was given in (4.3), $b_{2}$ is given by the fitted formula [76]
$b_{2}(z)=0.1802+0.1711z+0.02161z^{2}+0.03136z^{3}$ (4.14)
@@
×$\times[1+A_{0}(F(k_{1})+F(k_{2}))]+2\ \text{perm}\Big{]}$ (4.15)
@@
×$\times[1+g_{0}(g(k_{1})\mathcal{P}_{2}(\mu_{1})+g(k_{2})\mathcal{P}_{2}(\mu_{2}))]+2\text{perm}\Big{\}}$ (4.16)
In equations (4.15) and (4.2), we have assumed that both dipolar and quadrupolar asymmetry parameters are small so that only first-order terms $\mathcal{O}(A_{0})$ and $\mathcal{O}(g_{0})$ need to be considered.
@@ 5 Results and Discussion
In this section, we use Fisher formalism [56, 44, 78] to constrain dipolar ($A_{0}$) and quadrupolar ($g_{0}$) asymmetry parameters. As we have only one parameter in either of the cases, the Fisher matrix will have only one element.
Figure 2: Cumulative error $\sigma_{A_{0}}$ on the dipolar anisotropy parameter $A_{0}$ using 21cm power spectrum for SKA-Mid, HIRAX and PUMA surveys. The results are shown with the range relevant for the considered survey (see Table 1). We have shown our results for $k_{\mathrm{fg}}=0.01h\mathrm{Mpc}^{-1}$ and $k_{\mathrm{fg}}=0.02h\mathrm{Mpc}^{-1}$. Plots from left to right correspond to various scaling $F(k)=(1-k/k_{A})^{n}$, for $n=1,2,3$ and $k_{A}=1$ Mpc (see §3.1 for more details). For a given $z$ in the common range, PUMA gives the best constraints.
Figure 3: Cumulative error $\sigma_{g_{0}}$ on the quadrupolar anisotropy parameter $g_{0}$ using 21cm power spectrum for SKA-Mid, HIRAX and PUMA. The results are shown for various scaling $g(k)=(k/k_{*})^{n}$, discussed in §3.2. Top panel: From left to right, $n=-2,-1$ and bottom panel: $n=0,1,2$. Just like for dipolar anisotropy, in this case too, PUMA gives the best constraints in the common range.

### v2

@@ 1 Introduction
[…]
@@
[…]
@@ 2.3 Foreground Problem and its Avoidance
$|k_{\parallel}|>\tan\alpha(z)k_{\perp},\$(2.8)
@@ 3.1 Dipolar Asymmetry
[…]
@@
where $\hat{\mathbf{p}}$ is the direction of asymmetry, $\hat{\mathbf{n}}$ the vector along the LoS which we choose to be along Z axis for convenience and $P_{\zeta}^{0}(k)$ was defined above. Also, we have taken a scale dependent modulation $A(k)$, motivated by the fact that the hemispherical power asymmetry in CMB is present at large angular scales and typically dies out beyond $\ell\gtrsim 64$ [5]. The scale dependence is also suggested by quasar density, where at low $z$ the dipole modulation vanishes at scales $k\sim 1\ \mathrm{Mpc}^{-1}$ [12]. Parameterizing the scale dependent amplitude as $A(k)=A^{\prime}F(k)$, with $A^{\prime}$ denoting the amplitude, we can write (3.2) as
@@
where $A_{0}=2A^{\prime}\hat{p}\cdot\hat{n}=2A^{\prime}\hat{p}\cdot\hat{z}$ is the parameter we later constrain using Fisher formalism. Further, for the present study, we consider the following scale dependencies $F(k)$
$F(k)=\left(1-\frac{k}{k_{A}}\right)^{n},\ n=0,2$ (3.4)
where $k_{A}=1\ \mathrm{Mpc}^{-1}$. The case $n=0$ corresponds to no scale dependence that we compare with $n=2$ used in [22] to discuss the detectability of 21cm power spectrum. This is also motivated by some theoretical models where varying spectral index $n_{\mathrm{s}}$ can destroy asymmetry at some scale $k_{A}$ [66]. In addition to this, from CMB observations, a decaying behaviour is captured by the function
$F(k)=\left(\frac{k}{k^{c}_{A}}\right)^{-0.5}$ (3.5)
where $k_{A}^{c}=0.005\ \mathrm{Mpc}^{-1}$ [59, 67, 68]. We use all three parameterisations for our Fisher analysis results.
@@ 3.2 Quadrupolar Asymmetry
Although statistical isotropy is the prediction of standard inflationary theory, theoretical research has shown that inflationary models can be constructed that deviate from it. One such model of primordial anisotropy is when the inflaton interacts with a vector field having a nonzero vacuum expectation, transferring the anisotropy [17]. Such a model provides a quadrupolar anisotropic primordial power spectrum. We consider the primordial scalar power spectrum to have the following quadrupolar directional dependence about some axis $\hat{\mathbf{d}}$ [58, 25, 69]. In the coordinate system LoS is parallel to $\hat{z}$.
[…]
Then owing to the fact that the power spectrum is real, this expression can be written as
$P_{\zeta}(\mathbf{k})\approx P^{0}_{\zeta}\left[1+g(k)\sum_{i}g_{i}\mathfrak{F}_{i}(\mu,\phi)\right]$ (3.7)
real parameters [^1] Of course, there is no relation between $g_{i}$’s in (3.8) & $g_{0}$ and $g_{1}$ used in (2.1) that represented hyperfine level degeneracies used before. $g_{i}$ are chosen for ease of calculations, $\mu=\mathbf{\hat{k}}\cdot\hat{z}$ and $\phi$ is the polar angle of $\mathbf{k}$
$g_{1}=\sqrt{\frac{8\pi}{15}}g_{*}\Re(Y_{22}^{*}(\mathbf{\hat{d}})),\ g_{2}=-\sqrt{\frac{8\pi}{15}}g_{*}\Im(Y_{22}^{*}(\mathbf{\hat{d}})),\ {g_{3}=-\sqrt{\frac{32\pi}{15}}g_{*}\Re(Y_{21}^{*}(\mathbf{\hat{d}}))}$,
[…]
here $\Re(z)$ and $\Im(z)$ respectively denote the real and imaginary parts of a complex number $z$ and functions $\mathfrak{F}_{i}$ are defined as
$\mathfrak{F}_{1}=(1-\mu^{2})\cos 2\phi,\ \mathfrak{F}_{2}=(1-\mu^{2})\sin 2\phi,\ \mathfrak{F}_{3}=\mu\sqrt{1-\mu^{2}}\cos\phi,$
$\mathfrak{F}_{4}=\mu\sqrt{1-\mu^{2}}\sin\phi,\ \mathfrak{F}_{5}=3\mu^{2}-1$ (3.9)
@@
here $k_{*}=0.05\mathrm{Mpc}^{-1}$ is the pivot scale chosen by Planck [25]. We later constrain parameters $g_{i}$ using Fisher formalism.
@@ 4.1 Power Spectrum
where $\bar{T}_{\mathrm{b}}$ is the mean brightness temperature, the term $D_{\mathrm{FoG}}(\mathbf{k})=\text{exp}[-k^{2}\mu^{2}\sigma^{2}_{\mathrm{P}}(z)]$ corresponds to the Fingers of God (FoG) effect, with $\sigma_{\mathrm{P}}$ denoting the LoS dispersion in relative galaxy velocities. In this paper, we use $\sigma_{\mathrm{p}}=350/H(z)$, as per Ref. [73]. Further, $b_{1}$ is the neutral hydrogen bias given by [74]
$b_{1}(z)=0.754+0.0877z+0.0607z^{2}-0.00274z^{3}$ (4.3)
@@
where $\mathcal{T}(k)$ is the relevant transfer function and 0 in the superscripts are the quantities in the absence of primordial asymmetries. In the presence of asymmetries, this equation gets modified to
@@
$\mathrm{Quadrupolar}$ $:P_{\mathrm{m}}^{0}(k,z)\longrightarrow P_{\mathrm{m}}(\mathbf{k},z)=P_{\mathrm{m}}^{0}(k,z)\Big{[}1+g(k)\sum_{i}g_{i}\mathfrak{F}_{i}(\mu,\phi)\Big{]}$ (4.7)
where the scales dependencies $F(k)$ and $g(k)$ are given in (3.4) and (3.10) and $\mathfrak{F}$ are given in (3.9). Recently, it has been pointed out in [75, 76] that due to non-zero quadrupolar anisotropy, expression (4.7) should also contain a bias term, say $b_{1}^{(2)}$, corresponding to a tidal field at linear order. Ref [76] discusses $b_{\mathrm{h}}^{(2)}$ in this context while relating halo overdensity with dark matter density fluctuations. The authors estimate this parameter using simulations and find that the bias factor $b_{\mathrm{h}}^{(2)}$ is nearly scale-invariant on linear scale. We believe such a term is important for a more careful analysis but would require an estimation of the bias factor $b_{1}^{(2)}$. Due to present unavailability of such results, we neglect the contribution arising from this term for both power spectrum and bispectrum analysis.
Thus, our final expressions for the 21cm power spectrum in the presence of the dipolar and quadrupolar anisotropies are simplified to
[…]
$P_{21}^{\mathrm{Q}}(\mathbf{k},z)$ $=P^{0}_{21}(\mathbf{k},z)+g(k)\sum_{i}g_{i}\mathfrak{F}_{i}(\mu,\phi)(P^{0}_{21}(\mathbf{k},z)-P_{\mathrm{N}}(\mathbf{k},z))$ (4.9)
@@ 4.2 Bispectrum
The Dirac Delta imposes (a) the condition that the vectors $\mathbf{k}_{1}$, $\mathbf{k}_{2}$ and $\mathbf{k}_{3}$ form three sides of the triangle and (b) homogeneity as before. Apparently, there are 9 variables (three components each for $\mathbf{k}_{i}$, $i=1,2,3$) in the bispectrum expression (4.10) but three variables are eliminated by the Dirac Delta condition $\mathbf{k}_{1}+\mathbf{k}_{2}+\mathbf{k}_{3}=0$, thereby leading to 6 independent variables. Also, the bispectrum is invariant under a rotation in the plane perpendicular to the LoS [77]. So we are left with 5 variables that completely describe the bispectrum. We can choose three variables to define the shape of the triangle and remaining two for describing its orientation with respect to the LoS. Thus finally, 21cm bispectrum, (4.10) can be written as $B_{21}(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{k}_{3},z)=B_{21}(k_{1},k_{2},k_{3},\mu,\phi,z)$, where $k_{1}$, $k_{2}$ and $k_{3}$ are the sides of the triangle, $\mu$ is the cosine of angle between $\mathbf{k}_{1}$ and Z axis (the chosen LoS). Also, $\phi$ is the azimuthal angle about $\mathbf{k}_{1}$ [70] (see discussion in Appendix A).
@@
here $\mu_{12}=(\mu_{1}k_{1}+\mu_{2}k_{2})/k_{12}$, $k_{12}^{2}=(\mathbf{k}_{1}+\mathbf{k}_{2})^{2}$, $F_{2}(\mathbf{k}_{1},\mathbf{k}_{2})$ & $G_{2}(\mathbf{k}_{1},\mathbf{k}_{2})$ are the mode-coupling part of the matter density contrast & matter velocity [79] and $S_{2}(\mathbf{k}_{1},\mathbf{k}_{2})=(\hat{k}_{1}\cdot\hat{k}_{2})^{2}-1/3$ is the tidal kernel [80]. The neutral hydrogen biases $b_{i}(z)$, $i=1,2$ and $b_{s^{2}}$ relate the statistics of observed tracers with the dark matter distribution. The bias $b_{1}$ was given in (4.3), $b_{2}$ is given by the fitted formula [74]
$b_{2}(z)=-0.308-0.0724z-0.0534z^{2}+0.0247z^{3}$ (4.14)
@@
$\times[1+A_{0}(F(k_{1})+F(k_{2}))]+2\ \text{perm}\Big{]}$ (4.15)
@@
$\times\big{[}1+\sum_{i}g_{i}\{g(k_{1})\mathfrak{F}_{i}(\mu_{1},\phi_{1})+g(k_{2})\mathfrak{F}_{i}(\mu_{2},\phi_{2})\}\big{]}+2\text{perm}\Big{\}}$ (4.16)
In (4.15) and (4.16), we assume both dipolar and quadrupolar asymmetry parameters are small so that we only need to consider first-order terms $\mathcal{O}(A_{0})$ and $\mathcal{O}(g_{i})$. Also, the polar angles $\phi_{i}$ in (4.16) are calculated in Appendix A.
@@ 5 Results and Discussion
In this section, we use Fisher formalism [56, 44, 83] to constrain dipolar ($A_{0}$) and quadrupolar ($g_{i}$, $i=1,2,\ldots,5$) asymmetry parameters. As we’ve only one parameter in case of dipolar asymmetry, the Fisher matrix will have only one element for either power spectrum or bispectrum. For the quadrupolar case, the size of the matrix depends upon whether we evaluate power spectrum or bispectrum. As we discuss later, we find that in case of power spectrum, $g_{1}$ shows degeneracy with $g_{2}$ and $g_{3}$ with $g_{4}$. Thus effectively there are only 3 parameters to constrain when we consider power spectrum. We find no such simplification in case of bispectrum and need to constrain all 5 parameters.
Figure 2: Cumulative error $\sigma_{A_{0}}$ on the dipolar anisotropy parameter $A_{0}$ using 21cm power spectrum for SKA-Mid, HIRAX and PUMA surveys. The results are shown with the redshift range relevant for the considered survey (see Table 1). We have shown our results for both $k_{\mathrm{fg}}=0.01h\mathrm{Mpc}^{-1}$ and $0.02h\mathrm{Mpc}^{-1}$. First two plots correspond to the scaling $F(k)=(1-k/k_{A})^{n}$ for $n=0,2$ and $k_{A}=1\ \mathrm{Mpc}^{-1}$ while the rightmost one is for $F(k)=(k/k^{c}_{A})^{-0.5}$ with $k_{A}^{c}=0.005\mathrm{Mpc}^{-1}$ (see §3.1 for more details). For a given $z$ in the common range, PUMA gives the best constraints irrespective of the chosen scale dependence.
Figure 3: Cumulative error $\sigma_{g_{i}}$ on the quadrupolar anisotropy parameters $g_{i}$ using power spectrum for SKA, HIRAX and PUMA surveys (respectively from left to right). The results are shown with the range relevant for the considered survey (see Table 1) for both $k_{\mathrm{fg}}=0.01h\mathrm{Mpc}^{-1}$ and $0.02h\mathrm{Mpc}^{-1}$. We’ve chosen the scale dependence $g(k)=1$ in these plots. As can be seen, the best constraints are obtained on $g_{5}$.
Figure 4: Cumulative error $\sigma_{g_{5}}$ on the quadrupolar anisotropy parameter $g_{5}$ using 21cm power spectrum for SKA-Mid, HIRAX and PUMA. In these plots, we focus exclusively on $g_{5}$, we find to be the best constrained parameter. The results are shown for various scalings $g(k)=(k/k_{*})^{n}$, discussed in §3.2. Top panel: From left to right, n=2,1n=-2,-1 and bottom panel: $n=1,2$. Just like for dipolar anisotropy, in this case too, PUMA gives the best constraints in the common range.